# Interactive Multiscale Modeling to Bridge Atomic Properties and Electrochemical Performance in Li-CO2 Battery Design

### v5

@@ 3.3.1 Li-CO2 battery electrolyte transport properties
The cations transference number is assessed by [94]:
Where the self-diffusion coefficients are determined from the mean-square-displacement of center of mass presented in Figure 3 (b), calculated from the MD simulations as follows [95]:
In the expression above, represents the temporal position of a random walker (ion or CO2 molecule). Here, indicates the time and is an initial time at which the random walker begins to move. The diffusion coefficients are extrapolated from the normal diffusion regime characterized by:
where is crossover time to the normal diffusion regime.
@@
3.3.2 Li+ solvation structure from large scale MD simulation
The structural analysis was conducted utilizing the Radial Distribution Function (RDF), g(r), and Coordination Number (Nc). The RDF quantifies the spatial distribution of particles relative to a reference particle, providing insights into the local structural organization that affect the transport properties in the electrolyte.
@@ 3.3.2 Li+ solvation structure from large scale MD simulation
The primary peak in g(r) corresponds to the nearest neighbors, indicating the most probable interparticle distance and interaction strength. This relationship is described by , where and is the interaction potential. Subsequent peaks in the RDF elucidate higher-order neighbor relationships, offering insights into medium-range order, coordination numbers, and packing efficiency. In isotropic, homogeneous systems, g(r) asymptotically approaches unity at large distances, signifying uniform particle distribution.
@@
We present the average Radial Distribution Function () and Coordination Number () profiles for BF, TFSI-, and EMIM+ ions, as well as DMSO solvent, surrounding a single Li+ ion. Additionally, we analyze the and between molecules of the same type, represented by their center-of-mass atoms. Li+ ions preferentially coordinate with molecules containing highly electronegative atoms, i.e., atoms with a negative partial charge. The flat curves in Figure 3 (e) indicate a weak interaction between Li+ and EMIM+. Consequently, Li+ primarily coordinates with BF, TFSI-, and DMSO. According to Figure 3 (c), (d), and (f), Li+ ions interact strongly with BF, TFSI-, and DMSO. The most probable distances between Li+ and the constituent atoms are: BF (Li+-F 2.47 Å, Li+-B 3.22 Å), TFSI- (Li+-O 2.57 Å, Li+-F 2.85 Å), and DMSO (Li+-O 2.45 Å). The Li+ cations in the electrolyte are more solvated by BF and DMSO than TFSI-, as evidenced by the higher peaks in the and the corresponding coordination number curves of Li+-BF. BF and DMSO are closer to the Li+ ions than TFSI- ions, playing a crucial role in the dissociation of LiTFSI salt. However, TFSI- can also access the first coordination shell of Li+ due to strong interactions between its O and F atoms and Li+. The presence of both BF and TFSI- facilitates favorable coordination between Li+ and F, potentially leading to the formation of a Li-F solid electrolyte interphase (SEI) that protects the Li-metal anode. Interactions between molecules of the same type are weak due to electrostatic repulsion, resulting in well-dispersed molecules throughout the simulated electrolyte. The average first shell coordination number values for a tagged Li+ cation were determined using a threshold distance of 3 Å, based on the range of the first peaks. Key findings from the coordination number analysis include: Li+-Li+ (, ), Li+-BF (, ), Li+-DMSO (), and Li+-TFSI- (, , ).
@@
The high ionic conductivity and transference number (Table 2) of the simulated electrolyte is derived from several key factors revealed by the solvation structure and interaction analyses (Figure 3). Li+ cations are preferentially solvated by BF and DMSO, promoting LiTFSI salt dissociation and increasing free charge carriers. Weak Li+-TFSI- interactions facilitate easier Li+ transport. The presence of BF and TFSI- creates Li+ conducting channels through favorable Li+-F coordination. Well-dispersed ionic species due to weak same-type molecular interactions enhance efficient ion transport. Optimal coordination numbers (e.g., for BF, for DMSO) allow efficient Li+ transport while maintaining sufficient solvation. Low Li+-Li+ coordination () minimizes clustering, promoting ion mobility. Multiple solvating species create a dynamic environment, facilitating rapid solvating molecule exchange and enhancing Li+ mobility. The preferential solvation of Li+ by BF and DMSO allows for more free movement compared to the larger TFSI- anions, resulting in a higher proportion of the ionic current carried by Li+ and thus increasing the Li+ transference number.
@@ 3.4 Cell performance from finite element analysis
The simulated voltage-capacity curves of the Li-CO2 battery cell during charge/discharge cycles are shown in Figure 4(a). These curves were generated through finite element analysis (FEA) using input parameters derived from atomistic simulations (Figure 2, Table 2) and additional material properties detailed in Tables S1–S3 (Supporting Information). The FEA results were validated against experimental data collected at a current density of 1 mAh/cm2, as described in the Supporting Information (SI). A strong correlation between simulated and experimental voltages is observed. During discharge, the voltage profile exhibits three distinct regions: (1) an initial sharp voltage drop due to kinetic overpotential () [98], (2) a linear decline attributed to resistive losses () across the growing Li2CO3/C film (average plateau potential: 2.65 V), (3) a "sudden death" characterized by rapid voltage collapse after 5000 mAh/g. The latter two phenomena, previously described for Li-O2 batteries as charge-transport limitations through Li2O2 films [99], exhibit analogous behavior in Li-CO2 systems. During charging, the cell maintains a stable potential of approximately 4.2 V until reaching 6000 mAh/g, followed by a steep voltage rise. The system achieves a specific capacity of 6200 mAh/g (2.48 mAh/cm2) normalized to the catalyst mass.
@@
where: = 7500 mAh/g is the baseline capacity, = 110000 mAh/g is a scaling factor, = 0.157 is a characteristic current density. At low current density (i 0.157 ), the capacity tends towards its maximum value of 117500 mAh/g. This is the sum of the baseline capacity () and the additional capacity represented by . As the current density increases, there is a rapid decrease in the additional capacity provided by , resulting in a decrease in the total capacity . The parameter determines the rate at which the capacity decreases with increasing current density. A smaller would mean a faster decrease in capacity for a given increase in current density, while a larger would indicate a slower decrease. As the current density increases further, the capacity approaches the baseline value of 7500 mAh/g. This suggests that at high current densities, the battery’s capacity is primarily limited to its baseline capacity, as the contribution from the exponential term becomes negligible. The exponential decrease in capacity with increasing current density highlights a trade-off between power output and energy capacity [102]. At higher current densities (i 0.157 ), the battery can deliver more power, but at the cost of reduced capacity. This is a critical consideration in applications where both high-power and high-capacity are required. The connection between cell voltage () and current density () during the discharge of the studied Li-CO2 battery cell in Figure 4 (d) is represented by the formula [101, 103]:
@@
where V and V. The parameter represents the slope of the voltage change with the logarithm of the current density/(mA/cm2), while represents the voltage at a reference current density of 1 mA/cm2. This formula indicates that as the current density increases, the cell voltage decreases logarithmically, a common characteristic in electrochemical cells due to increased polarization and decreased voltage at higher current densities [101, 103]. The performance of Li-CO2 batteries is governed by a complex interplay of structural and operational parameters, including the initial cathode porosity, CO2 diffusion in the battery electrolyte, initial CO2 concentration in the cathode, and the kinetics of the cathodic reaction. These factors collectively influence key performance metrics, such as capacity and voltage plateau presented at a current density of 0.4 mA/cm2 in Figure S2 and Figure 4 (e) and (f). Our analysis reveals a strong positive correlation between porosity () and capacity (), as demonstrated by the best linear fit equation:
This trend is attributed to the increased surface area and improved CO2 accessibility in higher porosity structures, which facilitate more efficient electrochemical reactions. Similarly, the CO2 diffusion coefficient () and initial CO2 concentration () exhibit significant positive correlations with capacity, as shown by the best linear fit equations:
These results highlight the critical role of CO2 transport and availability in determining battery capacity. Enhanced diffusion and higher CO2 concentrations improve the accessibility of CO2 to active sites, thereby promoting more efficient electrochemical reactions. This is consistent with recent findings by Chen et al., who demonstrated that a Li-O2/CO2 battery can achieve a full discharge capacity of 6628 mAh/g and a long life of 715 cycles, surpassing the performance of pure Li-O2 batteries [104]. However, the cathodic reaction rate constant () shows a logarithmic relationship with both capacity and voltage plateau, as evidenced by:
This logarithmic relationship underscores the significant impact of reaction kinetics on battery performance, emphasizing the need for efficient catalysts to optimize reaction rates. Furthermore, both CO2 diffusion and concentration demonstrate positive correlations with the voltage plateau, as shown by:
These findings indicate that improving CO2 transport and availability not only enhances capacity but also stabilizes the voltage plateau, which is critical for achieving stable and efficient battery operation. A recent study demonstrated a high-performance Li-CO2 battery using a Cu(II) solid redox mediator, achieving a higher discharge voltage of 2.8 V and a lower charge potential of 3.7 V [25]. Therefore, optimizing the cathode porosity, CO2 diffusion and concentration, while improving reaction kinetics through advanced catalysts, are essential strategies for enhancing the performance of Li-CO2 batteries. Future research should focus on developing innovative electrode materials, electrolyte and catalytic systems to leverage these relationships and advance the practical application of Li-CO2 batteries as a sustainable energy storage technology. For instance, a recent study reported a high-rate Li-CO2 battery enabled by a 2D medium-entropy catalyst, which operates at a high current density of 5000 mA/g and capacity of 5000 mAh/g for up to 125 cycles, far exceeding previously reported values [32].
@@
Figure 5 (a1) and (a2) illustrate how porosity () and volume fraction of Li2CO3/C deposited species () change over space and time with current densities of i=1 mA/cm2 and i=0.1 mA/cm2, respectively. Figure S3 presents a comparison of these parameters at i=0.1, i=0.25, i=0.5, and i=1 mA/cm2. For high current density (i=1 mA/cm2), porosity decreases gradually, primarily near the CO2 feed side of the cathode (X/L=1), from its initial value of to 0.48 at the end of discharge. This decrease is due to the formation of a thin film of Li2CO3/C, represented by the gradual increase of from 0 to 0.25 near the cathode feed side. This film formation leads to electrode clogging, impeding efficient CO2 diffusion into the cathode. For low current density (i=0.1 mA/cm2), Li2CO3/C forms larger structures [105], evidenced by the gradual increase of across a wider region of the cathode from 0 to 0.64. This results in a distributed porosity decrease from 0.73 to 0.09, initially allowing more CO2 transport into the cathode. However, as discharge progresses, the substantial porosity decrease hinders CO2 diffusion even at low current densities. Figure 5 (b1) and (b2) present the CO2 concentration profiles within the cell during discharge at i=1 mA/cm2 and i=0.1 mA/cm2, respectively. Figure S4 compares concentration profiles at various current densities. At i=1 mA/cm2, the CO2 concentration () decreases sharply from its initial value of 4.8 mol/m3 to 0 mol/m3 in the porous cathode, indicating strongly impeded CO2 diffusion. At i=0.1 mA/cm2, the decrease is more gradual. The impediment to CO2 transport, which is inversely proportional to the current density, results from the accumulation of Li2CO3 and C on the active cathode surface. This accumulation reduces pore availability for the electrolyte, increases the amount of insulating species, and obstructs pores, collectively restricting CO2 transport and limiting cell capacity (Figure 4 (c)). The nucleation and growth of Li2CO3 significantly impact performance, as described in Figure 2 (c) and the detailed FEA model in section 2.4. Current density plays a crucial role in this process. At high current densities (1 mA/cm2), rapid nucleation near the CO2 feed side forms a thin Li2CO3/C film (Figure 5 (a1), (b1)). In contrast, low current densities (0.1 mA/cm2) promote slower growth, resulting in larger, more distributed Li2CO3 structures and a gradual decrease in porosity (Figure 5 (a2), (b2)).
To address these limitations, an ideal cathode should feature: large surface area for reaction sites and product deposition, large pores to facilitate active species transport and prevent clogging, high porosity to accommodate discharge product growth, and high electrical conductivity and stability. Such a cathode design could mitigate the observed limitations and potentially enhance the overall performance of Li-CO2 batteries [106].

### v6

@@ 3.3.1 Li-CO2 battery electrolyte transport properties
The cations’ transference number is assessed by [66]:
Where the self-diffusion coefficients are determined from the mean-square-displacement of the center of mass presented in Figure 5 (b), calculated from the MD simulations as follows [67]:
In the expression above, represents the temporal position of a random walker (ion or CO2 molecule). Here, indicates the time, and is the initial time at which the random walker begins to move. The diffusion coefficients are extrapolated from the normal diffusion regime characterized by:
where is the crossover time to the normal diffusion regime.
@@
The calculated diffusion coefficients, transference number, and ionic conductivity are directly incorporated into the FEA component of our multiscale modeling framework. These atomistically derived transport properties serve as critical input parameters for the FEA, enabling us to accurately simulate the intercalation and transport of CO2 and ions during battery charge/discharge cycling at the cell level.
3.3.2 Li+ solvation structure from large-scale MD simulation
The structural analysis was conducted utilizing the Radial Distribution Function (RDF), g(r), and Coordination Number (Nc). The RDF quantifies the spatial distribution of particles relative to a reference particle, providing insights into the local structural organization that affects the transport properties in the electrolyte.
@@ 3.3.2 Li+ solvation structure from large-scale MD simulation
The primary peak in g(r) corresponds to the nearest neighbors, indicating the most probable interparticle distance and interaction strength. This relationship is described by , where , and is the interaction potential. Subsequent peaks in the RDF elucidate higher-order neighbor relationships, offering insights into medium-range order, coordination numbers, and packing efficiency. In isotropic, homogeneous systems, g(r) asymptotically approaches unity at large distances, signifying a uniform particle distribution.
@@
We present the average Radial Distribution Function () and Coordination Number () profiles for BF, TFSI-, and EMIM+ ions, as well as DMSO solvent, surrounding a single Li+ ion. Additionally, we analyze the and between molecules of the same type, represented by their center-of-mass atoms. Li+ ions preferentially coordinate with molecules containing highly electronegative atoms, i.e., atoms with a negative partial charge. The flat curves in Figure 5 (e) indicate a weak interaction between Li+ and EMIM+. Consequently, Li+ primarily coordinates with BF, TFSI-, and DMSO. According to Figure 5 (c), (d), and (f), Li+ ions interact strongly with BF, TFSI-, and DMSO. The most probable distances between Li+ and the constituent atoms are: BF (Li+-F 2.47 Å, Li+-B 3.22 Å), TFSI- (Li+-O 2.57 Å, Li+-F 2.85 Å), and DMSO (Li+-O 2.45 Å). The Li+ cations in the electrolyte are more solvated by BF and DMSO than TFSI-, as evidenced by the higher peaks in the and the corresponding coordination number curves of Li+-BF. BF and DMSO are closer to the Li+ ions than TFSI- ions, playing a crucial role in the dissociation of LiTFSI salt. However, TFSI- can also access the first coordination shell of Li+ due to strong interactions between its O and F atoms and Li+. The presence of both BF and TFSI- facilitates favorable coordination between Li+ and F, potentially leading to the formation of a Li-F solid electrolyte interphase (SEI) that protects the Li-metal anode. Interactions between molecules of the same type are weak due to electrostatic repulsion, resulting in well-dispersed molecules throughout the simulated electrolyte. The average first shell coordination number values for a tagged Li+ cation were determined using a threshold distance of 3 Å, based on the range of the first peaks. Key findings from the coordination number analysis include: Li+-Li+ ) and ), Li+-BF (, ), Li+-DMSO (), and Li+-TFSI- ( and ).
@@
The high ionic conductivity (Table 5) of the simulated electrolyte is derived from several key factors revealed by the solvation structure and interaction analyses (Figure 5). Li+ cations are preferentially solvated by BF and DMSO, promoting LiTFSI salt dissociation and increasing free charge carriers. Weak Li+-TFSI- interactions facilitate easier Li+ transport. The presence of BF and TFSI- creates Li+ conducting channels through favorable Li+-F coordination. Well-dispersed ionic species due to weak same-type molecular interactions enhance efficient ion transport. Optimal coordination numbers (e.g., for BF, for DMSO) allow efficient Li+ transport while maintaining sufficient solvation. Low Li+-Li+ coordination () minimizes clustering, promoting ion mobility. Multiple solvating species create a dynamic environment, facilitating rapid solvating molecule exchange and enhancing Li+ mobility. The preferential solvation of Li+ by BF and DMSO allows for more free movement compared to the larger TFSI- anions, resulting in a higher proportion of the ionic current carried by Li+ and thus increasing the Li+ transference number.
@@ 3.4 Cell performance from Finite Element Analysis
The simulated voltage-capacity curves of the Li-CO2 battery cell during charge/discharge cycles are shown in Figure 6 (a). These curves were generated through Finite Element Analysis (FEA) using input parameters derived from atomistic simulations (Figure 3, Table 5) and additional material properties detailed in Tables 1, 2, and 3. The FEA results were validated against experimental data collected at a current density of 1 mAh/cm2, as described in the Supporting Information (SI). A strong correlation between simulated and experimental voltages is observed. During discharge, the voltage profile exhibits three distinct regions: (1) an initial sharp voltage drop due to kinetic overpotential () [70], (2) a linear decline attributed to resistive losses () across the growing Li2CO3/C film (average plateau potential: 2.65 V), (3) a "sudden death" characterized by rapid voltage collapse after 5000 mAh/g. The latter two phenomena, previously described for Li-O2 batteries as charge-transport limitations through Li2O2 films [71], exhibit analogous behavior in Li-CO2 systems. During charging, the cell maintains a stable potential of approximately 4.2 V until reaching 6000 mAh/g, followed by a steep voltage rise. The system achieves a specific capacity of 6200 mAh/g (0.48 mAh/cm2).
@@
where: = 7500 mAh/g is the baseline capacity, = 110000 mAh/g is a scaling factor, = 0.157 is a characteristic current density. At low current density (i 0.157 ), the capacity tends towards its maximum value of 117500 mAh/g. This is the sum of the baseline capacity () and the additional capacity represented by . As the current density increases, there is a rapid decrease in the additional capacity provided by , resulting in a decrease in the total capacity . The parameter determines the rate at which the capacity decreases with increasing current density. A smaller would mean a faster decrease in capacity for a given increase in current density, while a larger would indicate a slower decrease. As the current density increases further, the capacity approaches the baseline value of 7500 mAh/g. This suggests that at high current densities, the battery’s capacity is primarily limited to its baseline capacity, as the contribution from the exponential term becomes negligible. The exponential decrease in capacity with increasing current density highlights a trade-off between power output and energy capacity [74]. At higher current densities (i 0.157 ), the battery can deliver more power, but at the cost of reduced capacity. This is a critical consideration in applications where both high power and high capacity are required. The connection between cell voltage () and current density () during the discharge of the studied Li-CO2 battery cell in Figure 6 (d) is represented by the formula [73, 75]:
@@
where V and V. The parameter represents the slope of the voltage change with the logarithm of the current density/(mA/cm2), while represents the voltage at a reference current density of 1 mA/cm2. This formula indicates that as the current density increases, the cell voltage decreases logarithmically, a common characteristic in electrochemical cells due to increased polarization and decreased voltage at higher current densities [73, 75].
Figure 7 summarizes the influence of key parameters on the voltage–capacity profiles of Li-CO2 batteries at a current density of 1 mA/cm2. To investigate the influence of key parameters on the voltage-capacity profiles (Figure 7), we performed a parameter sensitivity analysis using FEA to solve the model equations described in Section 2.4. The base case parameters, as provided in Tables 1, 2, and 3, were used as the reference. Each parameter was varied over a specified range while keeping the others at their base values. Specifically, the initial porosity () was varied from 0.6 to 0.9 in steps of 0.05, the cathodic reaction rate constant () from 1 10-19 to 1 10-14 m7/s/mol2 (logarithmically spaced), the initial CO2 concentration (c) from 2.8 to 6.8 mol/m3, the cathode thickness () from 0.05 to 0.3 mm, the ionic conductivity () from 1 to 10 mS/cm, and the CO2 diffusion coefficient () from 0.6 10-9 m2/s to 1.6 10-9 m2/s. For each set of parameters, the discharge process was simulated at a constant current density of 1 mA/cm2. As shown in Figure 7 (a), increasing the initial porosity () leads to higher battery capacity by expanding the available surface area and enhancing CO2 accessibility, thereby promoting more efficient electrochemical reactions. Figure 7 (b) highlights the critical role of the cathodic reaction rate constant (); the voltage–capacity profile exhibits high sensitivity to variations in , emphasizing the necessity of effective catalytic materials for optimal battery performance. In Figure 7 (c), it is evident that higher initial CO2 concentration () results in increased capacity and a more stable voltage plateau, attributed to improved reactant availability at the cathode. Figure 7 (d) demonstrates that increasing the cathode thickness () from 0.05 mm to 0.1 mm substantially enhances capacity due to the greater active volume; however, further increases beyond 0.1 mm do not provide additional capacity gains, likely due to diffusion limitations. Figure 7 (e) shows that electrolyte ionic conductivity () is a critical factor, with values above 2 mS/cm required to achieve high capacities (>4000 mAh/g) and improved voltage stability. Figure 7 (f) indicates that a higher CO2 diffusion coefficient () further enhances both capacity and voltage plateau by facilitating more efficient CO2 transport to active sites. Notably, the same trends are observed at lower current densities; Figure S2 depicts these effects at a current density of 0.4 mA/cm2. Taken together, these results underscore the importance of simultaneously optimizing electrode structure, electrolyte properties, and catalytic activity to maximize the performance of Li-CO2 batteries.
@@
Figure 8 (a1) and (a2) illustrate how porosity () and volume fraction of Li2CO3/C deposited species () change over space and time with current densities of i=1 mA/cm2 and i=0.1 mA/cm2, respectively. Figure S3 presents a comparison of these parameters at i=0.1, i=0.25, i=0.5, and i=1 mA/cm2. For high current density (i=1 mA/cm2), porosity decreases gradually, primarily near the CO2 feed side of the cathode (X/L=1), from its initial value of to 0.48 at the end of discharge. This decrease is due to the formation of a thin film (flake-like) of Li2CO3/C [76], represented by the gradual increase of from 0 to 0.25 near the cathode feed side. This film formation leads to electrode clogging, impeding efficient CO2 diffusion into the cathode. For low current density (i=0.1 mA/cm2), Li2CO3/C forms larger structures [76], evidenced by the gradual increase of across a wider region of the cathode from 0 to 0.64. This results in a distributed porosity decrease from 0.73 to 0.09, initially allowing more CO2 transport into the cathode. However, as discharge progresses, the substantial porosity decrease hinders CO2 diffusion even at low current densities. Figure 8 (b1) and (b2) present the CO2 concentration profiles within the cell during discharge at i=1 mA/cm2 and i=0.1 mA/cm2, respectively. Figure S4 compares concentration profiles at various current densities. At i=1 mA/cm2, the CO2 concentration () decreases sharply from its initial value of 4.8 mol/m3 to 0 mol/m3 in the porous cathode, indicating strongly impeded CO2 diffusion. At i=0.1 mA/cm2, the decrease is more gradual. The impediment to CO2 transport, which is inversely proportional to the current density, results from the accumulation of Li2CO3 and C on the active cathode surface. This accumulation reduces pore availability for the electrolyte, increases the amount of insulating species, and obstructs pores, collectively restricting CO2 transport and limiting cell capacity (Figure 6 (c)). The nucleation and growth of Li2CO3 significantly impact performance, as described in Figure 4 and the detailed FEA model in section 2.4. Current density plays a crucial role in this process. At high current densities (1 mA/cm2), rapid nucleation near the CO2 feed side forms a thin Li2CO3/C film (Figure 8 (a1), (b1)). In contrast, low current densities (0.1 mA/cm2) promote slower growth, resulting in larger, more distributed Li2CO3 structures and a gradual decrease in porosity (Figure 8 (a2), (b2)).
To address these limitations, an ideal cathode should feature: a large surface area for reaction sites and product deposition, large pores to facilitate active species transport and prevent clogging, high porosity to accommodate discharge product growth, and high electrical conductivity and stability. Such a cathode design could mitigate the observed limitations and potentially enhance the overall performance of Li-CO2 batteries [77]. Finally, we note that direct experimental validation is essential to confirm our model’s mechanistic predictions. To rigorously assess accuracy and applicability, high-resolution imaging techniques such as transmission electron microscopy (TEM) and scanning electron microscopy (SEM) should be used to verify predicted changes in Li2CO3/C morphology, cathode surface features, and Li2CO3/C aggregation states. Additionally, time-resolved or ex-situ X-ray diffraction (XRD) can track the evolution of Li2CO3/C crystal structure and phase composition, while gas adsorption measurements, including Brunauer-Emmett-Teller (BET) and Barrett-Joyner-Halenda (BJH) analysis, can provide a quantitative assessment of cathode porosity and specific surface area. Importantly, a multiscale experimental approach is also needed to capture material behavior across different length and time scales, ensuring that both nanoscale mechanisms and macroscopic properties are accurately linked and described. These targeted and multiscale experiments are crucial for establishing the model’s predictive accuracy and real-world relevance.